One year ago, you purchased a 2-year bond with YTM of 5% and coupon rate of 4% (semiannual payments). When you are selling it today, you notice that its YTM has decreased to 3% and bond price has changed accordingly. You 1-year holding period return is roughly:
Answer #1

Assuming a face value of $1000

Purchase price:  Using financial calculator
Input: FV= 1000,PMT=4%*1000/2 = 20, I/Y=5/2= 2.5

N = 2*2 = 4

Solve for PV as -981.19

Hence P0=981.19

Selling price:

Using financial calculator
Input: FV= 1000,PMT=4%*1000/2 = 20, I/Y=3/2= 1.5

N = 1*2 = 2

Solve for PV as -1009.78

Hence P1= 1009.78

Holding period return = (P1-P0+Coupon)/P0

= (1009.78-981.19+20)/981.19

= 4.95%